OpenStudy (anonymous):

I GIVE MEDALS Evaluate the following and write the answer in scientific notation: (6 x 10^-5)/ (5 x 10^-8) A. 1.2 x 10^3 B. 1.2 x 10^-13 C. 1.2 x 10^-3

OpenStudy (anonymous):

SOMEONE PLEASE HELP

OpenStudy (ranga):

The Equation editor doesn't seem to be working. (6 x 10^-5)/ (5 x 10^-8) = 6 / 5 * 10^(-5) / 10^(-8) = 1.2 X 10^(-5 + 8) = 1.2 X 10^3